# Beyond boundaries: Gary Lorden’s groundbreaking contributions to sequential analysis

### v1

@@ Abstract
Gary Lorden provided a number of fundamental and novel insights to sequential hypothesis testing and changepoint detection. In this article we provide an overview of Lorden’s contributions in the context of existing results in those areas, and some extensions made possible by Lorden’s work, mentioning also areas of application including threat detection in physical-computer systems, near-Earth space informatics, epidemiology, clinical trials, and finance.
@@
Sequential hypothesis testing; changepoint detection; CUSUM; multihypothesis sequential tests.
@@ 1 Introduction
The purpose of this article is to provide an overview of Gary Lorden’s contributions to the field of sequential analysis. Lorden obtained fundamental and wide-reaching results in sequential hypothesis testing and changepoint detection, which we aim to present in the context of those areas. Beginning with hypothesis testing in Section 2, after describing the testing setup and optimality of the sequential probability ratio test (SPRT, Section 2.2), we cover Lorden’s fundamental inequality for excess over the boundary (Section 2.3), Lorden’s results on multi-parameter testing and their application to near-optimality of the multihypothesis SPRT (Section 2.4), Lorden’s contributions to the Keifer-Weiss problem of testing while minimizing the expected sample size at a parameter value between the hypotheses and other results building off of his (Section 2.5), optimal testing of composite hypotheses (Section 2.6), and optimal multistage testing (Section 2.7). In Section 3 we cover Lorden’s fundamental minimax changepoint detection theory and related results in the field. Finally, in Section 4, we mention some of Lorden’s consulting work, both within and outside academia.
@@ 2 Sequential Hypothesis Testing
2.1 General Formulation of the Multihypothesis Testing Problem
@@
$p_{i,n}({\mathbf{X}}^{n})=\prod_{t=1}^{n}f_{i,t}(X_{t}|{\mathbf{X}}^{t-1}),\quad i=0,1,\dots,N,$ (1)
where $f_{i,n}(X_{n}|{\mathbf{X}}^{n-1})$, $n\geqslant 1$ are corresponding conditional densities.
For $n\in\mathbb{N}$, define the likelihood ratio (LR) process between the hypotheses $\operatorname{\mathcal{H}}_{i}$ and $\operatorname{\mathcal{H}}_{j}$ as
$\Lambda_{ij}(n)=\frac{{\mathrm{d}}{\mathsf{P}}_{i}^{n}}{{\mathrm{d}}{\mathsf{P}}_{j}^{n}}({\mathbf{X}}^{n})=\frac{p_{i,n}({\mathbf{X}}^{n})}{p_{j,n}({\mathbf{X}}^{n})}=\prod_{t=1}^{n}\frac{f_{i,t}(X_{t}|{\mathbf{X}}^{t-1})}{f_{j,t}(X_{t}|{\mathbf{X}}^{t-1})}$
and the log-likelihood ratio (LLR) process as
$\lambda_{ij}(n)=\log\Lambda_{ij}(n)=\sum_{t=1}^{n}\log\left[\frac{f_{i,t}(X_{t}|{\mathbf{X}}^{t-1})}{f_{j,t}(X_{t}|{\mathbf{X}}^{t-1})}\right].$
In the particular case of i.i.d. observations, under hypothesis $\operatorname{\mathcal{H}}_{i}$ the observations are independent and identically distributed with a common density $f_{i}(X_{n})$, $n=1,2,\dots$, so the joint density $p_{i,n}({\mathbf{X}}^{n})=p_{i}({\mathbf{X}}^{n})$ in (1) is
$p_{i}({\mathbf{X}}^{n})=\prod_{t=1}^{n}f_{i}(X_{t}),\quad i=0,1,\dots,N.$ (2)
A multihypothesis sequential test is a pair $\delta=(T,d)$, where $T$ is a stopping time with respect to the filtration $\{{\mathscr{F}}_{n}\}_{n\in\mathbb{Z}_{+}}$ and $d=d({\mathbf{X}}^{T})$ is an ${\mathscr{F}}_{T}$-measurable terminal decision function with values in the set $\{0,1,\dots,N\}$. Specifically, $d=i$ means that the hypothesis $\operatorname{\mathcal{H}}_{i}$ is accepted upon stopping, $\left\{d=i\right\}=\left\{T<\infty,~{}\delta~{}\text{accepts~{}$\operatorname{\mathcal{H}}_{i}$}\right\}$. Let $\alpha_{ij}(\delta)={\mathsf{P}}_{i}(d=j)$, $i\neq j$, $i,j=0,1,\dots,N$, denote the error probabilities of the test $\delta$, i.e., the probabilities of accepting the hypothesis $\operatorname{\mathcal{H}}_{j}$ when $\operatorname{\mathcal{H}}_{i}$ is true.
@@
${\mathbb{C}}({\bm{\alpha}})=\left\{\delta:\alpha_{ij}(\delta)\leqslant\alpha_{ij}~{}\text{for}~{}i,j=0,1,\dots,N,\,i\neq j\right\},$ (3)
where ${\bm{\alpha}}=(\alpha_{ij})$ is a matrix of given error probabilities that are positive numbers less than $1$.
Let ${\mathsf{E}}_{i}$ denote the expectation under the hypothesis $\operatorname{\mathcal{H}}_{i}$ (i.e., under the measure ${\mathsf{P}}_{i}$). The goal of a statistician is to find a sequential test that would minimize the expected sample sizes ${\mathsf{E}}_{i}[T]$ for all hypotheses $\operatorname{\mathcal{H}}_{i}$, $i=0,1,\dots,N$ at least approximately, say asymptotically for small probabilities of errors as $\alpha_{ij}\to 0$.
2.2 Optimality of Wald’s SPRT for Testing Two Hypotheses
Assume first that $N=1$, i.e., that we are dealing with two hypotheses $\operatorname{\mathcal{H}}_{0}$ and $\operatorname{\mathcal{H}}_{1}$. In the mid 1940s, Wald (1945, 1947) introduced the Sequential Probability Ratio Test (SPRT) for the sequence of i.i.d. observations $X_{1},X_{2},\dots$, in which case the LR $\Lambda_{1,0}(n)=\Lambda_{n}$ is $\Lambda_{n}=\prod_{t=1}^{n}~{}[f_{1}(X_{t})/f_{0}(X_{t})]$. After $n$ observations have been made Wald’s SPRT prescribes for each $n\geqslant 1$:
stop and $\text{stop and accept}~{}~{}\operatorname{\mathcal{H}}_{1}$ $\text{if}\quad\Lambda_{n}\geqslant A_{1};$
stop and $\text{stop and accept}~{}~{}\operatorname{\mathcal{H}}_{0}$ $\text{if}\quad\Lambda_{n}\leqslant A_{0};$
continue sampling $\text{if}\quad A_{0}<\Lambda_{n}<A_{1},$
where $A_{0}<1<A_{1}$ are two thresholds.
Let $Z_{t}=\log[f_{1}(X_{t})/f_{0}(X_{t})]$ be the LLR for the observation $X_{t}$, so the LLR for the sample ${\mathbf{X}}^{n}$ is the sum
$\lambda_{10}(n)=\lambda_{n}=\sum_{t=1}^{n}Z_{t},\quad n=1,2,\dots$ (4)
Let $a_{0}=\log A_{0}<0$ and $a_{1}=\log A_{1}>0$. The SPRT $\delta_{*}(a_{0},a_{1})=(T_{*},d_{*})$ can be represented in the form
$T_{*}(a_{0},a_{1})=\inf\left\{n\geqslant 1:\lambda_{n}\notin(a_{0},a_{1})\right\},\quad d_{*}(a_{0},a_{1})=\begin{cases}1&\text{if}~{}~{}\lambda_{T_{*}}\geqslant a_{1}\\ 0&\text{if}~{}~{}\lambda_{T_{*}}\leqslant a_{0}.\end{cases}$ (5)
In the case of two hypotheses, the class of tests (3) is of the form
@@ 2.2 Optimality of Wald’s SPRT for Testing Two Hypotheses
i.e., it upper-bounds the probabilities of errors of Type 1 (false positive) $\alpha_{0}(\delta)=\alpha_{0,1}(\delta)$ and Type 2 (false negative) $\alpha_{1}(\delta)=\alpha_{1,0}(\delta)$, respectively.
@@
Wald’s SPRT has an extraordinary optimality property: it minimizes both expected sample sizes ${\mathsf{E}}_{0}[T]$ and ${\mathsf{E}}_{1}[T]$ in the class of sequential (and non-sequential) tests ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ with given error probabilities as long as the observations are i.i.d. under both hypotheses. More specifically, Wald and Wolfowitz (1948) proved, using a Bayesian approach, that if $\alpha_{0}+\alpha_{1}<1$ and thresholds $a_{0}$ and $a_{1}$ can be selected in such a way that $\alpha_{0}(\delta_{*})=\alpha_{0}$ and $\alpha_{1}(\delta_{*})=\alpha_{1}$, then the SPRT $\delta_{*}$ is strictly optimal in the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$, i.e.,
${\mathsf{E}}_{i}[T_{*}]=\inf_{\delta\in{\mathbb{C}}(\alpha_{0},\alpha_{1})}{\mathsf{E}}_{i}[T]\quad\text{for}~{}i=0,1.$
@@
Lai (1981) proved that the SPRT is also first-order asymptotically optimal for general non-i.i.d. models (1) with dependent and non-identically distributed observations when the normalized log-likelihood ratio $n^{-1}\lambda_{n}$ converges $1$-quickly to finite numbers $I_{i}$ under ${\mathsf{P}}_{i}$, $i=0,1$, that is

### v2

@@ Abstract
Gary Lorden provided several fundamental and novel insights into sequential hypothesis testing and changepoint detection. In this article, we provide an overview of Lorden’s contributions in the context of existing results in those areas, and some extensions made possible by Lorden’s work. We also mention some of Lorden’s significant consulting work, including as an expert witness and for NASA, the entertainment industry, and Major League Baseball.
@@
Sequential hypothesis testing; changepoint detection; CUSUM; multihypothesis sequential tests; Caltech.
@@ 1 Introduction
The purpose of this article is to provide an overview of Gary Lorden’s significant contributions to the field of sequential analysis. But first, to give a sense of Lorden’s remarkable life and personality, in Sections 1.1-1.2 we give a brief biography of Lorden and highlight some of his extra-academic work.
Figure 1: Counter-clockwise from top: Gary Lorden as an undergraduate at Caltech in 1959; lecturing in 1987; and lecturing circa 2010. Photos courtesy of Caltech.
1.1 Biographical Sketch
Gary Allen Lorden was born in Los Angeles, California, on June 10, 1941. Lorden entered Caltech as a freshman in 1958, and Lorden’s undergraduate contemporaries at Caltech included a number of others who would also go on to be notable statisticians including Larry Brown, Peter Bickel, Brad Efron, and Carl Morris. Lorden received a BS from Caltech in 1962 and a PhD from Cornell University in 1966 under the supervision of Jack Kiefer. After a faculty position at Northwestern University, Lorden rejoined Caltech in 1968, where he stayed until his retirement in 2009.
Beyond research and teaching, Lorden was known for his leadership roles at Caltech. He served as dean of students from 1984 to 1988, vice president for student affairs from 1989 to 1998, and acting vice president for student affairs in 2002. He was executive officer (Caltech’s version of department chair) for the mathematics department from 2003 to 2006.
Gary and his wife, Louise, were both accomplished pianists and enjoyed playing and singing duets for guests, often students, at their home in Pasadena. They also enjoyed showing students how well they danced together. Lorden also liked to act and regularly participated in plays put on at Caltech.
Lorden passed away on October 25, 2023 at the age of 82. He is survived by his children Lisa and Diana, and his wife Louise passed away in 2015.
1.2 Consulting, Hollywood, and Major League Baseball
Lorden was known for his creativity and generosity of ideas, so it is not surprising that throughout his career he was in demand as an academic collaborator and consultant, and much of his early work on changepoint detection arose from collaborations with researchers at the Jet Propulsion Laboratory (JPL) in Pasadena, California, which is managed for NASA by Lorden’s home institution of Caltech.
But Lorden was also an uncommonly effective, engaging, and entertaining communicator and this caused many outside academics to seek him out as a consultant too. Lorden would routinely be interviewed by reporters and appear on the evening news, explaining things like the odds of winning the latest lottery jackpot. Lorden also “moonlighted” as an expert witness in court cases involving statistics and math.
Eventually Hollywood came calling, and in 2005 Lorden was asked to be the math consultant for a (then) new CBS television show called NUMB3RS. In a bit of art imitating life, the show was about a Caltech professor who used math to help the FBI solve crimes, and Lorden worked with the show’s writers to accurately incorporate mathematical topics into the storylines. An aspect of the relationship that Lorden particulalry relished, that he related to us, was that the show’s creators initially considered setting the show at MIT (Caltech’s rival), but decided to change the venue after learning more about Caltech and Lorden. The show would go on to be a hit, running for 118 episodes over 6 seasons. If a viewer of the show knew something of Lorden’s work they could see some of his favorite topics (and many topics discussed below in this article) woven into the episodes’ plot lines including changepoint detection, hypothesis testing, Bayesian methods, gambling math, cryptography, and sports statistics among others. With Keith Devlin, Lorden wrote a popular general audience book (Devlin and Lorden 2007) on the mathematical topics appearing in the show. For example, Devlin and Lorden (2007, Chapter 4) explain the basic concept of changepoint detection:
and goes on to discuss this method’s importance for quick response to potential bioterrorist attacks and for designing efficient algorithms to pinpoint various kinds of criminal activity, such as to detect an increase in crime rates in certain geographical areas and to track changes in financial transactions that could be criminal.
1.3 The Remainder of this Article
@@
The remainder of this article is devoted to describing Lorden’s fundamental and far-reaching results in sequential hypothesis testing and changepoint detection, which we aim to present within the context of those areas. Beginning with hypothesis testing in Section 2, after describing the testing setup and optimality of the sequential probability ratio test (SPRT) in Section 2.1.1, Lorden’s findings on multi-parameter testing and their application to near-optimality of the multihypothesis SPRT are covered in Section 2.1.2. We then cover Lorden’s fundamental inequality for excess over the boundary in Section 2.2. Additionally, we explore Lorden’s contributions to the Keifer-Weiss problem of testing while minimizing the expected sample size at a parameter value between the hypotheses and other results stemming from his work (Section 2.3), optimal testing of composite hypotheses (Section 2.4), and optimal multistage testing (Section 2.5). In Section 3 we cover Lorden’s fundamental minimax changepoint detection theory and related advancements in the field.
@@ 2 Sequential Hypothesis Testing
In this section, we delve into Lorden’s contributions to hypothesis testing, encompassing the challenges of excess over the boundaries of random walks, the formulation of nearly optimal multi-decision sequential rules (including multihypothesis tests and multistage tests), and the modified Keifer–Weiss problem. These topics were extensively explored in the seminal Lorden’s papers (Lorden 1967, 1970, 1976, 1977a, 1972, 1973, 1980, 1983).
2.1 Multihypothesis Testing
2.1.1 The General Multihypothesis Testing Problem
One of Lorden’s major fundamental contributions is the proposal of a multihypothesis sequential test that achieves third-order asymptotic optimality in the i.i.d. case, akin to Wald’s SPRT when error probabilities are small. Further details will be provided in Subsection 2.1.2. Additionally, extensions to a more general non-i.i.d. case, where observations may exhibit dependency and non-identical distributions, have been explored in Lai (1981); Tartakovsky (1998, 2020, 2024); Tartakovsky, Nikiforov, and Basseville (2015).
@@ 2.1.1 The General Multihypothesis Testing Problem
Assume that the observed random variables $X_{1},X_{2},\dots$ are independent and identically distributed (i.i.d.), so that under ${\mathsf{P}}_{i}$ the sample ${\mathbf{X}}^{n}=(X_{1},\dots,X_{n})$ has joint density $p_{i}({\mathbf{X}}^{n})$ with respect to the dominating measure ${\mathsf{Q}}^{n}$ for all $n\in\mathbb{N}$, which can be expressed as
$p_{i}({\mathbf{X}}^{n})=\prod_{t=1}^{n}f_{i}(X_{t}),\quad i=1,\dots,N,$ (1)
where $f_{i}(X_{t})$ represents the respective density for $X_{t}$ under the hypothesis $\operatorname{\mathcal{H}}_{i}$. Hence, the interest lies in determining which of $N$ given densities $f_{1},\dots,f_{N}$ is true.
A multihypothesis sequential test is a pair $\delta=(T,d)$, where $T$ is a stopping time with respect to the filtration $\{{\mathscr{F}}_{n}\}_{n\in\mathbb{Z}_{+}}$ and $d=d({\mathbf{X}}^{T})$ is an ${\mathscr{F}}_{T}$-measurable terminal decision function with values in the set $\{1,\dots,N\}$. Specifically, $d=i$ means that the hypothesis $\operatorname{\mathcal{H}}_{i}$ is accepted upon stopping, $\left\{d=i\right\}=\left\{T<\infty,~{}\delta~{}\text{accepts~{}$\operatorname{\mathcal{H}}_{i}$}\right\}$. Let $\alpha_{ij}(\delta)={\mathsf{P}}_{i}(d=j)$, $i\neq j$, $i,j=1,\dots,N$, denote the error probabilities of the test $\delta$, i.e., the probabilities of accepting the hypothesis $\operatorname{\mathcal{H}}_{j}$ when $\operatorname{\mathcal{H}}_{i}$ is true.
@@
${\mathbb{C}}({\bm{\alpha}})=\left\{\delta:\alpha_{ij}(\delta)\leqslant\alpha_{ij}~{}\text{for}~{}i,j=1,\dots,N,\,i\neq j\right\},$ (2)
where ${\bm{\alpha}}=(\alpha_{ij})$ is a matrix of given error probabilities that are positive numbers less than $1$ (the diagonal entries $\alpha_{ii}$ are immaterial).
Let ${\mathsf{E}}_{i}$ denote the expectation under the hypothesis $\operatorname{\mathcal{H}}_{i}$ (i.e., under the measure ${\mathsf{P}}_{i}$). The objective is to discover a sequential test that would minimize the expected sample sizes ${\mathsf{E}}_{i}[T]$ for all hypotheses $\operatorname{\mathcal{H}}_{i}$, $i=1,\dots,N$, at least approximately.
For $n\in\mathbb{N}$, define the likelihood ratio (LR) and the log-likelihood ratio (LLR) processes between the hypotheses $\operatorname{\mathcal{H}}_{i}$ and $\operatorname{\mathcal{H}}_{j}$ as
$\Lambda_{ij}(n)$ $=\frac{{\mathrm{d}}{\mathsf{P}}_{i}^{n}}{{\mathrm{d}}{\mathsf{P}}_{j}^{n}}({\mathbf{X}}^{n})=\frac{p_{i}({\mathbf{X}}^{n})}{p_{j}({\mathbf{X}}^{n})}=\prod_{t=1}^{n}\frac{f_{i}(X_{t})}{f_{j}(X_{t})},$
$\lambda_{ij}(n)$ $=\log\Lambda_{ij}(n)=\sum_{t=1}^{n}\log\left[\frac{f_{i}(X_{t})}{f_{j}(X_{t})}\right].$
In a particular case of two hypotheses $\operatorname{\mathcal{H}}_{1}$ and $\operatorname{\mathcal{H}}_{2}$ ($N=2$), Wald (1945, 1947) introduced the Sequential Probability Ratio Test (SPRT). Let $Z_{t}=\log[f_{1}(X_{t})/f_{2}(X_{t})]$ be the LLR for the observation $X_{t}$, so the LLR for the sample ${\mathbf{X}}^{n}$ is the sum
$\lambda_{12}(n)=\lambda_{n}=\sum_{t=1}^{n}Z_{t},\quad n=1,2,\dots$ (3)
Letting $a_{0}<0$ and $a_{1}>0$ be thresholds, Wald’s SPRT $\delta_{*}(a_{0},a_{1})=(T_{*},d_{*})$ is
$T_{*}(a_{0},a_{1})=\inf\left\{n\geqslant 1:\lambda_{n}\notin(a_{0},a_{1})\right\},\quad d_{*}(a_{0},a_{1})=\begin{cases}1&\text{if}~{}~{}\lambda_{T_{*}}\geqslant a_{1}\\ 2&\text{if}~{}~{}\lambda_{T_{*}}\leqslant a_{0}.\end{cases}$ (4)
In the case of two hypotheses, the class of tests (2) is defined as
@@
where $\alpha_{0}$ represents the upper bound on the Type I error (false positive) probability $\alpha_{0}(\delta)=\alpha_{12}(\delta)$, and $\alpha_{1}$ represents the upper bound on the Type II error (false negative) probability $\alpha_{1}(\delta)=\alpha_{21}(\delta)$.
@@
Wald’s SPRT possesses an extraordinary optimality property: it minimizes both the expected sample sizes ${\mathsf{E}}_{1}[T]$ and ${\mathsf{E}}_{2}[T]$ within the class of sequential (and non-sequential) tests ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ as long as the observations are i.i.d. under both hypotheses. In other words: ${\mathsf{E}}_{i}[T_{*}]=\inf_{\delta\in{\mathbb{C}}(\alpha_{0},\alpha_{1})}{\mathsf{E}}_{i}[T]$ for $i=0,1$, as established by Wald and Wolfowitz (1948) through a Bayesian approach.
@@
Lai (1981) proved that the SPRT is also first-order asymptotically optimal as $\max(\alpha_{0},\alpha_{1})\to 0$ for general non-i.i.d. models with dependent and non-identically distributed observations when the normalized log-likelihood ratio $n^{-1}\lambda_{n}$ converges $1$-quickly to finite numbers $I_{i}$ under ${\mathsf{P}}_{i}$.
The central idea of Lorden’s investigation, elaborated in detail in Section 2.1.2, is that, similar to how the SPRT is strictly optimal in the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ for any error probabilities $\alpha_{0}$ and $\alpha_{1}$, combinations of SPRTs exhibit third-order asymptotic optimality for multihypothesis testing problems involving any finite number of densities when probabilities of errors are small.
2.1.2 Near Optimality of the Multihypothesis SPRT
The problem of sequentially testing many hypotheses is substantially more complex than that of testing two hypotheses. Identifying an optimal test in the class (2) that minimizes expected sample sizes for all hypotheses $\operatorname{\mathcal{H}}_{1},\dots,\operatorname{\mathcal{H}}_{N}$, is daunting. Hence, a significant portion of the development of sequential multihypothesis testing in the 20th century has focused on the exploration of certain combinations of one-sided SPRTs. See Armitage (1950); Chernoff (1959); Kiefer and Sacks (1963); Lorden (1967, 1977a).
The results of Lorden’s ingenious paper Lorden (1977a) are of fundamental importance as they establish third-order asymptotic optimality of the accepting multihypothesis test that he proposed. More specifically, Lorden established that just as the SPRT is optimal in the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ for testing two hypotheses, certain combinations of one-sided SPRTs are nearly optimal in a third-order sense in the class ${\mathbb{C}}({\bm{\alpha}})$, i.e., subject to error probability constraints expected sample sizes are minimized to within the negligible additive $o(1)$ term:
$\inf_{\delta\in{\mathbb{C}}({\bm{\alpha}})}{\mathsf{E}}_{i}[T]={\mathsf{E}}_{i}[T_{*}]+o(1)\quad\text{as}~{}{\alpha_{\rm max}}\to 0\quad\text{for all}~{}i=1,\dots,N,$ (5)
where ${\alpha_{\rm max}}=\max_{1\leqslant i,j\leqslant N,i\neq j}\alpha_{ij}$ and $T_{*}$ is the stopping time of the multihypothesis test $\delta_{*}$, which is defined below.
We now define a test proposed by Lorden, which we will refer to as the accepting Matrix SPRT. Write ${\mathscr{N}}=\{1,\dots,N\}$. For a threshold matrix $\mathbf{A}=(A_{ij})_{i,j\in{\mathscr{N}}}$, with $A_{ij}>0$ for $i\neq j$ and the $A_{ii}$ are immaterial (00, say), define the Matrix SPRT (MSPRT) $\delta_{*}=(T_{*},d_{*})$, built on one-sided SPRTs between the hypotheses $\operatorname{\mathcal{H}}_{i}$ and $\operatorname{\mathcal{H}}_{j}$, as follows:
and accept the unique $\operatorname{\mathcal{H}}_{i}$ that satisfies these inequalities. Note that for $N=2$ the MSPRT coincides with Wald’s SPRT.
$T_{i}=\inf\left\{n\geqslant 1:\min_{\stackrel{{\scriptstyle 1\leqslant j\leqslant N}}{{j\neq i}}}\left[\lambda_{ij}(n)-a_{ji}\right]\geqslant 0\right\},\quad i=1,\dots,N,$ (7)
the test in (6) can also be written in the following form:
$T_{*}=\min_{1\leqslant j\leqslant N}T_{j},\qquad d_{*}=i\quad\mbox{if}\quad T_{*}=T_{i}.$ (8)
Thus, in the MSPRT, each component SPRT is extended until, for some $i\in{\mathscr{N}}$, all $N-1$ SPRTs involving $\operatorname{\mathcal{H}}_{i}$ accept $\operatorname{\mathcal{H}}_{i}$.
The MSPRT is not strictly optimal for $N>2$ but it is a good approximation to the optimal multihypothesis test. Under certain conditions and with some choice of the threshold matrix $\mathbf{A}$, it minimizes the expected sample sizes ${\mathsf{E}}_{i}[T]$ for all $i=1,\dots,N$ to within a vanishing $o(1)$ term for small error probabilities; see (5).